x + 14y= 84
2x - 7y = -7
how do i solve these equations using substation?

Answer by checkley77(12844)

x + 14y= 84
x=84-14y substitute (84-14y) for x in the other equation & solve for y.
2x - 7y = -7
2(84-14)-7y=-7
168-28y-7y=-7
-35y=-7-168
-35y=-175
y=-175/-35
y=5 ans.
x=84-14*5
x=84-70
x=14 ans.
Proof:
2*14-7*5=-7
28-35=-7
-7=-7